OpenStudy (anonymous):

The graph of which equation has a vertex of (–3, 4)?

OpenStudy (anonymous):

a) y = 2(x – 3)2 + 4 b) y = 2(x + 3)2 + 4 c) y = 2(x – 4)2 + 3 d) y = 2(x + 4)2 – 3